OpenStudy (anonymous):

Find the equation of the line containing the point P and having the slope m. Write your answer in slope-intercept form. P = (-4, -6), m = 3

OpenStudy (aroub):

slope-intercept form = y=mx+b all you have to do is just substitute to find the y-intercept & form an equation y=mx+b , ( -4,6) m=3 6=3(-4)+b 6=-12 +b 18=b (m) slope= 3 & (b) y-intercept= 18 y=mx+b (substitute again...) => y=3x+18

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

so would y = 21

OpenStudy (aroub):

Why 21?

OpenStudy (anonymous):

u would add 3 + 18

OpenStudy (aroub):

No..no they are not like terms! you have 3x and 18 so y would just be 18 :)

OpenStudy (anonymous):

so the equation is y=3x+18

OpenStudy (aroub):

Yup!

OpenStudy (anonymous):

to graph the slope i put the point at 18 for y

OpenStudy (aroub):

(0,18)

OpenStudy (anonymous):

ok thanks thats where i was getting confused

OpenStudy (aroub):

You're welcome :)